Claim: Fewer than 94​% of adults have a cell phone. In a reputable poll of 1179 ​adults, 89​% said that they have a cell phone. Find the value of the test statistic.
Answer #1

Solution :

This is a left tailed test .

The null and alternative hypothesis is

H0 : p =0.94

Ha : p < 0.94

= 0.89

n =1179

P0 = 0.94

1 - P0 =1- 0.94 = 0.06

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

=0.89- 0.94 / [0.94*(1-0.94) /1179 ]

= -7.229

P(z < -7.229) = 0

P-value = 0